# Spinning masters: on the impact of tidal forces and protohalo size on early spin evolution

### v1

@@ 3.2 The tidal field at different scales
One of the key ingredients of the TTT is the tidal field associated with the position of each protohalo during the linear regime. However, it is commonly stated that, in order to eliminate the contribution of the protohalo itself to the perturbations responsible for its torquing, the potential field must be properly smoothed at protohalo scales before computing its second derivatives. This ensures not only that the resulting torque is external but also that the velocity flow is laminar. Although different methods have been adopted to perform this analysis, it is not clear which scale is most appropriate to correctly implement the TTT. In general, it is assumed that this scale is of the order of $R_{\mathrm{lag}}$, but applying a Gaussian or top-hat filter of this characteristic size in Fourier space eliminates not only the contribution of each protohalo but also that of all perturbations associated with similar or smaller wavelengths.
@@
In order to analyze the dependence of the TTT predictions on the characteristic scale of the tidal field, we carry out a systematic study of the effect that different smoothing lengths, $R_{\mathrm{s}}$, have on the tidal tensor $\mathbf{T}(R_{\mathrm{s}})=T_{ij}(R_{\mathrm{s}})$, through the following method. First, we determine the overdensity field $\delta(\mathbf{x})$ on a cubic grid of $1024^{3}$ cells, by means of a cloud in cell (CIC) interpolation. We then use NEXUS+ to compute, over this grid, the tidal tensor $T_{ij}(\mathbf{x})=\partial^{2}\Phi(\mathbf{x})/\partial x_{i}\partial x_{j}$, where $\Phi(\mathbf{x})=\phi(\mathbf{x})/(4\pi G\bar{\rho})$ and $\bar{\rho}$ is the mean comoving background density, by inverting in Fourier space the Poisson equation $\nabla^{2}\Phi=\delta(\mathbf{x})$. In this process we apply a series of Gaussian filters on the density field of the form $G(R_{\mathrm{s}})=e^{-k^{2}R_{\mathrm{s}}^{2}/2}$, where $R_{\mathrm{s}}$ takes $30$ logarithmically equispaced values between $R_{\mathrm{s}}=0.5{\textrm{ $h^{-1}$Mpc}}$ and 30 $30{\textrm{ $h^{-1}$Mpc}}$. This allows to produce multiple smoothed versions of the tidal field, $T_{ij}(\mathbf{x},R_{\mathrm{s}})$. We then interpolate these fields in coordinate space to the location $\mathbf{x}_{\mathrm{h}}$ of each halo. In this way we obtain a halo-centric catalogue of tidal tensor estimates with different smoothings, $T_{ij}(\mathbf{x}_{\mathrm{h}},R_{\mathrm{s}})$. Finally, in order to determine the preferred expansion or collapse directions, we diagonalize each tidal tensor and compute the corresponding eigenvalues and eigenvectors, $\lambda_{k}(\mathbf{x}_{\mathrm{h}},R_{\mathrm{s}})$ and $\hat{\mathbf{t}}_{k}(\mathbf{x}_{\mathrm{h}},R_{\mathrm{s}})$, with $k=1,2,3$, respectively. In what follows we will always assume that the eigenvalues and eigenvectors correspond to the halocentric tidal field, and therefore we will omit the dependence on the position of each proto-halo, $\mathbf{x}_{\mathrm{h}}$.
@@
Figure 3 illustrates the result of this procedure. Each row corresponds to the density field in the vicinity of a protohalo at redshift $z=80$. From top to bottom, the protohaloes have masses $M_{\mathrm{FoF}}=8.1\times 10^{11}$, $5.7\times 10^{12}$ and $1.9\times 10^{13}{\textrm{ $h^{-1}$M${}_{\sun}$}}$, respectively. In each case, the first column shows the particle distribution of the protohalo along with three different measurements of its size, $R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$, with dotted, continuous and dashed circles, respectively. The other three columns columns show the surrounding density field smoothed with gaussian kernels with radii $Rs/({\textrm{ $h^{-1}$Mpc}})=0.5,2.1$ and $8.4$, from left to right. The direction of the arrows indicates the orientation of the first and third axes of the smoothed tidal field, $\hat{\mathbf{t}}_{1}(R_{\mathrm{s}})$ and $\hat{\mathbf{t}}_{3}(R_{\mathrm{s}})$, respectively. The sign of the corresponding eigenvalues, $\lambda_{1}(R_{\mathrm{s}})$ and $\lambda_{3}(R_{\mathrm{s}})$, determines whether the matter is collapsing or expanding in that direction, whereas their absolute value is proportional to the length of the arrows.
@@ 4.1 Properties of the tidal field at $z=80$
In Figure 4 we present the distribution of $\log(\delta+1)$ for all protohaloes (upper row) and the variation of the median of each distribution as a function of protohalo mass (bottom row). The left-hand panels correspond to the results when computing the tidal tensor using $30$ equally log-spaced smoothing scales $R_{\mathrm{s}}$ between 0.5 $0.5{\textrm{ $h^{-1}$Mpc}}$ and 30 $30{\textrm{ $h^{-1}$Mpc}}$. Each value of $R_{\mathrm{s}}$ is represented by a coloured curve. With no surprise, the PDFs show that the expected value of $\delta$ typically approaches zero as $R_{\mathrm{s}}$ increases, while the dispersion narrows down. In other words, the overdensity associated with the tidal field becomes smaller when we consider longer wavelength modes. This is true for the general trend of the colored curves, but the specific values of $\delta$ for a given smoothing scale varies significantly around protohalos of different mass, especially when we consider small values of $R_{\mathrm{s}}$. This can be seen in the bottom-left panel: for $R_{\mathrm{s}}\sim 1{\textrm{ $h^{-1}$Mpc}}$, for example, the overdensity varies from $\log(\delta+1)\sim 0.006$ in the low mass extreme to $\sim 0.026$ for high mass protohalos. In general, a fixed smoothing scale yields higher values of $\delta$ for more massive protohalos, except for $R_{\mathrm{s}}<0.8{\textrm{ $h^{-1}$Mpc}}$ and $_{\mathrm{FoF}}>10^{13}{\textrm{ $h^{-1}$M${}_{\sun}$}}$ M. Conversely, for any fixed mass, the tendency is that the overdensity increases as $R_{\mathrm{s}}$ decreases.
@@
Figure 5: Distribution and median of the tidal shear, $\log(q^{2})$, which quantifies the anisotropy of the smoothed tidal fields. The top row shows the PDFs for fixed smoothing scales $R_{\mathrm{s}}$ (left) and for protohalo size-dependent filters corresponding to $R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$ (right). The bottom row shows the median values of $\log(q^{2})$ as a function of protohalo mass, $M_{\mathrm{FoF}}$, for the same filter scales. Error bars in the bottom-right panel indicate the interquartile range for three different mass bins.
@@
So far, the behaviour of the tidal field smoothed in different scales is as expected. What happens when, for each protohalo, we consider a value of $R_{\mathrm{s}}$ that takes into account its size? This is shown in the right-hand panels of Figure 4, where we present with black dotted, continuous and dashed lines the resulting overdensities when we compute the tidal tensor setting, for each protohalo, $R_{\mathrm{s}}=R_{\mathrm{eq}},R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$, respectively. When looking at the PDFs, we observe that the equivalent radius typically yields the higher overdensities, which is a consequence of $R_{\mathrm{eq}}$ being the smaller measure of protohalo size. However, the most interesting feature of this analysis can be seen in the relation between $\log(\delta+1)$ and the protohalo mass: the median of every distribution becomes practically mass-independent. In effect, the black curves in the bottom-right panel present little or no variation at all when we compare their value across the entire mass range of protohalos. As already noted, $R_{\mathrm{eq}}$ produces the stronger overdensities ($0.017\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$, $0.017\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$ and $0.016\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$, from left to right in the position of the error bars), whereas $R_{\mathrm{lag}}$ ($0.005\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$, $0.005\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$ and $0.006\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$) and $10R_{\mathrm{vir}}$ ($0.003\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$, $0.003\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$ and $0.002\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$) yield more homogeneous density distributions.
@@
In Figure 5, we present the distribution of the tidal shear. As in the previous figure, the upper panels show the PDF of $\log(q^{2})$ for all protohalos and different smoothing scales, whereas the bottom panels display the variation of the median of each distribution as a function of protohalo mass. Focusing on the top-left panel, we observe that the tidal field becomes, in general, more isotropic as larger values of $R_{\mathrm{s}}$ are chosen, i.e., there is a systematic shift toward lower values of $q^{2}$. This behavior results from the Gaussian filter suppressing the shorter-wavelength perturbations of the primordial density field. When we take into account the protohalo mass (bottom-left panel), the simplest trend is observed for large values of $R_{\mathrm{s}}$. At any fixed smoothing length beyond $\sim 4{\textrm{ $h^{-1}$Mpc}}$, the median of $\log(q^{2})$ becomes independent of the protohalo mass, showing that this scale represents somewhat of an upper bound for the correlation between the tidal anisotropy and the local overdensities. For values of $R_{\mathrm{s}}$ below $\sim 4{\textrm{ $h^{-1}$Mpc}}$, however, the slope of each coloured curve changes significantly with $M_{\mathrm{FoF}}$. In particular, for each smoothing scale there is a range of values of $M_{\mathrm{FoF}}$ within which the median of $\log(q^{2})$ descends abruptly. For $R_{\mathrm{s}}<1{\textrm{ $h^{-1}$Mpc}}$ the curves reach a minimum and then start increasing. The mass where this change of slope occurs is higher for larger values of $R_{\mathrm{s}}$.
@@
Finally, contrary to what we found for $\delta$, none of the protohalo size measurements produce a smoothed tidal field with a $q^{2}$ distribution that is independent of mass. When we consider their size as a filtering scale, protohalos of higher mass are typically embedded in more isotropic tidal fields, whether $R_{\mathrm{s}}$ is fixed to $R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$, or $10R_{\mathrm{vir}}$. For example, the median of $\log(q^{2})$ corresponding to $R_{\mathrm{s}}=R_{\mathrm{eq}}$ changes from $-3.1\raisebox{2.15277pt}{\tiny${}^{+0.2}_{-0.2}$}$ for protohalos in the low-mass subsample, to $-3.3\raisebox{2.15277pt}{\tiny${}^{+0.2}_{-0.2}$}$ for intermediate-mass protohalos, and $-3.5\raisebox{2.15277pt}{\tiny${}^{+0.2}_{-0.2}$}$ for high-mass protohalos.
@@ 5 Difference between true and predicted AM in the initial conditions
Interestingly, this relationship can be approximately accounted for when setting $R_{\mathrm{s}}=R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$, and $10R_{\mathrm{vir}}$. These results are shown in the right-hand panels of Figure 7. Surprisingly, using the Lagrangian radius $R_{\mathrm{lag}}$ does not yield the most accurate predictions when applying Eq. (4). Except for the lowest-mass protohalos, this scale tends to underestimate the AM amplitude, with the underestimation increasing for higher masses. For $R_{\mathrm{s}}=R_{\mathrm{lag}}$, the median of $\log[J_{\mathrm{TTT}}(R_{\mathrm{s}})/J(z_{80})]$ at the positions of the error bars is $-0.04\raisebox{2.15277pt}{\tiny${}^{+0.17}_{-0.13}$}$, $-0.11\raisebox{2.15277pt}{\tiny${}^{+0.19}_{-0.17}$}$, and $-0.14\raisebox{2.15277pt}{\tiny${}^{+0.17}_{-0.17}$}$, from left to right. A similar trend with slightly greater scatter is observed for $R_{\mathrm{s}}=10R_{\mathrm{vir}}$.
The protohalo size measure that yields the best prediction for the AM amplitude is the equivalent radius, $R_{\mathrm{eq}}$. The distribution for $R_{\mathrm{s}}=R_{\mathrm{eq}}$ in the top-right panel peaks near $J_{\mathrm{TTT}}(R_{\mathrm{s}})/J(z_{80})=1$ and is narrower than those for the other two scales. Additionally, the variation of the median ratio with protohalo mass (bottom-right panel)Z is less pronounced. Specifically, at the positions of the error bars, the medians are $0.04\raisebox{2.15277pt}{\tiny${}^{+0.17}_{-0.13}$}$, $0.03\raisebox{2.15277pt}{\tiny${}^{+0.13}_{-0.11}$}$, and $0.02\raisebox{2.15277pt}{\tiny${}^{+0.11}_{-0.10}$}$, from left to right.
@@
Figure 9: Evolution of the specific AM of protohalos, normalized to the expected growth factor predicted by the TTT, $a^{2}(t)\dot{D}(t)$. The solid gray curve in each panel represents the median evolution of $\log[j/a^{2}\dot{D}]$ for the full FoF halos, with the shaded region indicating the interquartile range. The dashed and dot-dashed gray curves show the same quantity but for particles that, at present time, reside within $0.1$ and $0.5$ of $R_{\mathrm{vir}}$, respectively. The dotted black line shows the prediction from the TTT (Eq. 4), where both the inertia and tidal tensors are computed at $z=80$ using a smoothing scale equal to the equivalent radius, $R_{\mathrm{eq}}$, of each protohalo. The shaded region at the right-hand side of each panel indicates the highly non-linear regime, where TTT predictions are not expected to hold. Each panel corresponds to protohalo samples of increasing mass, from left to right.
@@
  • The amplitude of the predicted AM varies significantly with both the smoothing scale $R_{\mathrm{s}}$ and protohalo mass $M_{\mathrm{FoF}}$. The most accurate prediction for the AM amplitude is obtained when the smoothing scale is set to the equivalent radius of each protohalo, $R_{\mathrm{eq}}$.

### v2

@@ 3.2 The tidal field at different scales
One of the key ingredients of the TTT is the tidal field associated with the position of each protohalo during the linear regime. However, it is commonly stated that, in order to eliminate the contribution of the protohalo itself to the perturbations responsible for its torquing, the potential field must be properly smoothed at protohalo scales before computing its second derivatives. This ensures not only that the resulting torque is external but also that the velocity flow is laminar. Although different methods have been adopted to perform this analysis, it is not clear which scale is most appropriate to correctly implement the TTT. In general, it is assumed that this scale is of the order of $R_{\mathrm{lag}}$, but applying a Gaussian or top-hat filter of this characteristic size in Fourier space suppresses not only the contribution from substructures and fluctuations within each protohalo, but also that of all perturbations associated with similar or smaller wavelengths.
@@
In order to analyze the dependence of the TTT predictions on the characteristic scale of the tidal field, we carry out a systematic study of the effect that different smoothing lengths, $R_{\mathrm{s}}$, have on the tidal tensor $\mathbf{T}(R_{\mathrm{s}})=T_{ij}(R_{\mathrm{s}})$, through the following method. First, we determine the overdensity field $\delta(\mathbf{x})$ on a cubic grid of $1024^{3}$ cells, by means of a cloud in cell (CIC) interpolation. We then use NEXUS+ to compute, over this grid, the tidal tensor $T_{ij}(\mathbf{x})=\partial^{2}\Phi(\mathbf{x})/\partial x_{i}\partial x_{j}$, where $\Phi(\mathbf{x})=\phi(\mathbf{x})/(4\pi G\bar{\rho})$ and $\bar{\rho}$ is the mean comoving background density, by inverting in Fourier space the Poisson equation $\nabla^{2}\Phi=\delta(\mathbf{x})$. In this process we apply a series of Gaussian filters on the density field of the form $G(R_{\mathrm{s}})=e^{-k^{2}R_{\mathrm{s}}^{2}/2}$, where $R_{\mathrm{s}}$ takes $30$ logarithmically equispaced values between $R_{\mathrm{s}}=0.5{\textrm{ $h^{-1}$Mpc}}$ and $30{\textrm{ $h^{-1}$Mpc}}$. This allows to produce multiple smoothed versions of the tidal field, $T_{ij}(\mathbf{x},R_{\mathrm{s}})$. We then interpolate these fields in coordinate space to the location $\mathbf{x}_{\mathrm{h}}$ of each halo. In this way we obtain a halo-centric catalogue of tidal tensor estimates with different smoothings, $T_{ij}(\mathbf{x}_{\mathrm{h}},R_{\mathrm{s}})$. Finally, in order to determine the preferred expansion or collapse directions, we diagonalize each tidal tensor and compute the corresponding eigenvalues and eigenvectors, $\lambda_{k}(\mathbf{x}_{\mathrm{h}},R_{\mathrm{s}})$ and $\hat{\mathbf{t}}_{k}(\mathbf{x}_{\mathrm{h}},R_{\mathrm{s}})$, with $k=1,2,3$, respectively. In what follows we will always assume that the eigenvalues and eigenvectors correspond to the halocentric tidal field, and therefore we will omit the dependence on the position of each proto-halo, $\mathbf{x}_{\mathrm{h}}$.
@@
Figure 3 illustrates the result of this procedure. Each row corresponds to the density field in the vicinity of a protohalo at redshift $z=80$. From top to bottom, the protohalos have masses $M_{\mathrm{FoF}}=8.1\times 10^{11}$, $5.7\times 10^{12}$ and $1.9\times 10^{13}{\textrm{ $h^{-1}$M${}_{\sun}$}}$, respectively. In each case, the first column shows the particle distribution of the protohalo along with three different measurements of its size, $R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$, with dotted, continuous and dashed circles, respectively. The other three columns columns show the surrounding density field smoothed with gaussian kernels with radii $Rs/({\textrm{ $h^{-1}$Mpc}})=0.5,2.1$ and $8.4$, from left to right. The direction of the arrows indicates the orientation of the first and third axes of the smoothed tidal field, $\hat{\mathbf{t}}_{1}(R_{\mathrm{s}})$ and $\hat{\mathbf{t}}_{3}(R_{\mathrm{s}})$, respectively. The sign of the corresponding eigenvalues, $\lambda_{1}(R_{\mathrm{s}})$ and $\lambda_{3}(R_{\mathrm{s}})$, determines whether the matter is collapsing or expanding in that direction, whereas their absolute value is proportional to the length of the arrows.
@@ 4.1 Properties of the tidal field at $z=80$
In Figure 4 we present the distribution of $\log(\delta+1)$ for all protohalos (upper row) and the variation of the median of each distribution as a function of protohalo mass (bottom row). The left-hand panels correspond to the results when computing the tidal tensor using $30$ equally log-spaced smoothing scales $R_{\mathrm{s}}$ between $0.5{\textrm{ $h^{-1}$Mpc}}$ and $30{\textrm{ $h^{-1}$Mpc}}$. Each value of $R_{\mathrm{s}}$ is represented by a coloured curve. With no surprise, the PDFs show that the expected value of $\delta$ typically approaches zero as $R_{\mathrm{s}}$ increases, while the dispersion narrows down. In other words, the overdensity associated with the tidal field becomes smaller when we consider longer wavelength modes. This is true for the general trend of the colored curves, but the specific values of $\delta$ for a given smoothing scale varies significantly around protohalos of different mass, especially when we consider small values of $R_{\mathrm{s}}$. This can be seen in the bottom-left panel: for $R_{\mathrm{s}}\sim 1{\textrm{ $h^{-1}$Mpc}}$, for example, the overdensity varies from $\log(\delta+1)\sim 0.006$ in the low mass extreme to $\sim 0.026$ for high mass protohalos. In general, a fixed smoothing scale yields higher values of $\delta$ for more massive protohalos, except for $R_{\mathrm{s}}<0.8{\textrm{ $h^{-1}$Mpc}}$ and $M_{\mathrm{FoF}}>10^{13}{\textrm{ $h^{-1}$M${}_{\sun}$}}$. Conversely, for any fixed mass, the tendency is that the overdensity increases as $R_{\mathrm{s}}$ decreases.
@@
Figure 5: Distribution and median of the tidal shear, $\log(q^{2})$, which quantifies the anisotropy of the smoothed tidal fields. The top row shows the PDFs for fixed smoothing scales $R_{\mathrm{s}}$ (left) and for protohalo size-dependent filters corresponding to $R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$ (right). The bottom row shows the median values of $\log(q^{2})$ as a function of protohalo mass, $M_{\mathrm{FoF}}$, for the same filter scales. Error bars in the bottom-right panel indicate the interquartile range for three different mass bins. In the bottom panels, curves corresponding to the highest $R_{\mathrm{s}}$ are omitted, as they show nearly constant values with negligible mass dependence and would obscure the behavior observed in the low-$R_{\mathrm{s}}$ curves.
@@
So far, the behaviour of the tidal field smoothed in different scales is as expected. What happens when, for each protohalo, we consider a value of $R_{\mathrm{s}}$ that takes into account its size? This is shown in the right-hand panels of Figure 4, where we present with black dotted, continuous and dashed lines the resulting overdensities when we compute the tidal tensor setting, for each protohalo, $R_{\mathrm{s}}=R_{\mathrm{eq}},R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$, respectively. When looking at the PDFs, we observe that the equivalent radius typically yields the higher overdensities, which is a consequence of $R_{\mathrm{eq}}$ being the smaller measure of protohalo size. However, the most interesting feature of this analysis can be seen in the relation between $\log(\delta+1)$ and the protohalo mass: the median of every distribution becomes practically mass-independent. In effect, the black curves in the bottom-right panel present little or no variation at all when we compare their value across the entire mass range of protohalos. As already noted, $R_{\mathrm{eq}}$ produces the stronger overdensities ($017\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}, 0.0$$017\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$ and $0.016\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$, from left to right in the position of the error bars), whereas $R_{\mathrm{lag}}$ ($005\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}, 0.0$$005\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$ and $0.006\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$) and $10R_{\mathrm{vir}}$ ($\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}, 0.0030.003$$\raisebox{2.15277pt}{\tiny${}^{+0.003}_{-0.003}$}$ and $0.002\raisebox{2.15277pt}{\tiny${}^{+0.002}_{-0.002}$}$) yield more homogeneous density distributions.
@@
In Figure 5, we present the distribution of $q^{2}$, i.e. the tidal shear. As in the previous figure, the upper panels show the PDF of $\log(q^{2})$ for all protohalos and different smoothing scales, whereas the bottom panels display the variation of the median of each distribution as a function of protohalo mass. Focusing on the top-left panel, we observe that the tidal field becomes, in general, more isotropic as larger values of $R_{\mathrm{s}}$ are chosen, i.e., there is a systematic shift toward lower values of $q^{2}$. This behavior results from the Gaussian filter suppressing the shorter-wavelength perturbations of the primordial density field. When we take into account the protohalo mass (bottom-left panel), the simplest trend is observed for large values of $R_{\mathrm{s}}$. At any fixed smoothing length beyond $\sim 4{\textrm{ $h^{-1}$Mpc}}$, the median of $\log(q^{2})$ becomes independent of the protohalo mass, showing that this scale represents somewhat of an upper bound for the correlation between the tidal anisotropy and the local overdensities. For values of $R_{\mathrm{s}}$ below $\sim 4{\textrm{ $h^{-1}$Mpc}}$, however, the slope of each coloured curve changes significantly with $M_{\mathrm{FoF}}$. In particular, for each smoothing scale there is a range of values of $M_{\mathrm{FoF}}$ within which the median of $\log(q^{2})$ descends abruptly. For $R_{\mathrm{s}}<1{\textrm{ $h^{-1}$Mpc}}$ the curves reach a minimum and then start increasing. The mass where this change of slope occurs is higher for larger values of $R_{\mathrm{s}}$.
@@
Finally, contrary to what we found for $\delta$, none of the protohalo size measurements produce a smoothed tidal field with a $q^{2}$ distribution that is independent of mass. When we consider their size as a filtering scale, protohalos of higher mass are typically embedded in more isotropic tidal fields, whether $R_{\mathrm{s}}$ is fixed to $R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$, or $10R_{\mathrm{vir}}$. For example, the median of $\log(q^{2})$ corresponding to $R_{\mathrm{s}}=R_{\mathrm{eq}}$ changes from $3.1\raisebox{2.15277pt}{\tiny${}^{+0.2}_{-0.2}$}$ for protohalos in the low-mass subsample, to $3.3\raisebox{2.15277pt}{\tiny${}^{+0.2}_{-0.2}$}$ for intermediate-mass protohalos, and $3.5\raisebox{2.15277pt}{\tiny${}^{+0.2}_{-0.2}$}$ for high-mass protohalos.
@@ 5 Difference between true and predicted AM in the initial conditions
Interestingly, this relationship can be approximately accounted for when setting $R_{\mathrm{s}}=R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$, and $10R_{\mathrm{vir}}$. These results are shown in the right-hand panels of Figure 7. Surprisingly, using the Lagrangian radius $R_{\mathrm{lag}}$ does not yield the most accurate predictions when applying Eq. (4). Except for the lowest-mass protohalos, this scale tends to underestimate the AM amplitude, with the underestimation increasing for higher masses. For $R_{\mathrm{s}}=R_{\mathrm{lag}}$, the median of $\log[J_{\mathrm{TTT}}(R_{\mathrm{s}})/J(z_{80})]$ at the positions of the error bars is $0.04\raisebox{2.15277pt}{\tiny${}^{+0.17}_{-0.13}$}$, $0.11\raisebox{2.15277pt}{\tiny${}^{+0.19}_{-0.17}$}$, and $0.14\raisebox{2.15277pt}{\tiny${}^{+0.17}_{-0.17}$}$, from left to right. A similar trend with slightly greater scatter is observed for $R_{\mathrm{s}}=10R_{\mathrm{vir}}$.
The protohalo size measure that yields the best prediction for the AM amplitude is the equivalent radius, $R_{\mathrm{eq}}$. The distribution for $R_{\mathrm{s}}=R_{\mathrm{eq}}$ in the top-right panel peaks near $J_{\mathrm{TTT}}(R_{\mathrm{s}})/J(z_{80})=1$ and is narrower than those for the other two scales. Additionally, the variation of the median ratio with protohalo mass (bottom-right panel) is less pronounced. Specifically, at the positions of the error bars, the medians are $04\raisebox{2.15277pt}{\tiny${}^{+0.17}_{-0.13}$}, 0.0$$03\raisebox{2.15277pt}{\tiny${}^{+0.13}_{-0.11}$}$, and $0.02\raisebox{2.15277pt}{\tiny${}^{+0.11}_{-0.10}$}$, from left to right.
@@
Figure 9: Evolution of the specific AM of protohalos, normalized to the expected growth factor predicted by the TTT, $a^{2}(t)\dot{D}(t)$. The solid gray curve in each panel represents the median evolution of $\log[j/a^{2}\dot{D}]$ for the full FoF halos, with the shaded region indicating the interquartile range. The dashed and dot-dashed gray curves show the same quantity but for particles that, at present time, reside within $0.1$ and $0.5$ of $R_{\mathrm{vir}}$, respectively. The thick dotted black line shows the prediction from the TTT (Eq. 4), where both the inertia and tidal tensors are computed at $z=80$ using a smoothing scale equal to the equivalent radius, $R_{\mathrm{eq}}$, of each protohalo. The thin dotted black line in the upper right corner represents the slope associated with constant angular momentum. The shaded region at the right-hand side of each panel indicates the highly non-linear regime, where TTT predictions are not expected to hold. Each panel corresponds to protohalo samples of increasing mass, from left to right.
@@
  • The amplitude of the predicted AM varies significantly with both the smoothing scale $R_{\mathrm{s}}$ and protohalo mass $M_{\mathrm{FoF}}$. Among the three size measures we tested ($R_{\mathrm{eq}}$, $R_{\mathrm{lag}}$ and $10R_{\mathrm{vir}}$, the most accurate prediction for the AM amplitude is obtained when the smoothing scale is set to the equivalent radius of each protohalo, $R_{\mathrm{eq}}$.